What mass of oxygen will react with 7.75g of P4?

1
Expert's answer
2022-04-29T10:36:04-0400

Let's write down the reaction equation:

"P_4 + 5O_2 => 2P_2O_5"

That means that 1 mole of P4 reacts with 5 moles of O2

Find the number of moles of P4:

"\nu_1 = m/M = 7.75/(31*4) = 0.0625 (mol)"

Find the number of moles of O2:

"\nu_2 = \nu_1*5 = 0.0625*5 = 0.3125 (mol)"

Find the mass of O2:

"m = M*\nu_2 = 32 * 0.3125 = 10 (g)"

Answer: 10g of oxygen will react with 7.75g of P4